DIFFERENTIAL GEOMETRY/PDE SEMINAR

Wednesday, May 12, 2004 Thomson 202 4:00 pm (Note: special day, time and room!)

Critical points and supersymmetric vacua

Michael Douglas (Rutgers University)

Supersymmetric vacua (‘universes’) of string/M theory may be identified with certain critical points of a holomorphic section (the ‘superpotential’) of a Hermitian holomorphic line bundle over a complex manifold. An important physical problem is to determine how many vacua there are and how they are distributed. The present paper initiates the study of the statistics of critical points ∇s = 0 of Gaussian random holomorphic sections with respect to a connection ∇. Even the expected number of critical points depends on the curvature of ∇. The principal results give formulas for the expected density and number of critical points of Gaussian random sections relative to ∇ in a variety of settings. The results are particularly concrete for Riemann surfaces. Analogous results on the density of critical points of fixed Morse index are given.

Based on joint work with Bernard Shiffman and Steve Zelditch.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.